Flora11:

the ratio of the number of men to the number of women in a club is 5:7. There are 12 more women than men. How many members are there altogether?

QuestionCoveBot:

Issue?

Flora11:

How should I start solving this question?

silvernight269:

First make it into an equation.

silvernight269:

We know that the population of women is 12 more than men. So its W = 12 + x

silvernight269:

From the ratio we can obviously tel that when added that is the total number of the population of men and women. But since its simplified attach X to both numbers. Making it 5x:7X. Add those two numbers. 5x+7x=12x

ramen:

you could do the way @silvernight269 did it or you could do it the manual way.

silvernight269:

What?

silvernight269:

Could you not tag me nxt time? thx

ramen:

k sorry

ramen:

since we know that men to women is 5:7 and there are two more women then men in the ratio, we could keep adding 5 to men and 7 to women until they have a difference of 12. This method is still within the ratio. So basically, we need a difference of 12 and bc there are 2 more women then men, 12/2 = 6. 6 would be the multiplier.

ramen:

we would end up with 5*6 and 7*6. i can't give direct answers so I'll leave you with that. but when multiplying the ratio by 6, you get a difference of 12 and all you need to do is add it all up for your answer.

Flora11:

Thank you everyone, that was very helpful!

ramen:

np

ramen:

did you get your answer?